OpenStudy (anonymous):

Professor Hines is performing a nuclear experiment in a particle accelerator. The particle has an initial position of -24 meters, an initial velocity of 1 m/s, and is subjected to a constant acceleration of 292 m/s/s for 1.1 seconds. All motion is in a straight line. ***I found the velocity, now I need the average velocity: What is the average velocity at the end of the time interval?

OpenStudy (tkhunny):

Have you considered the formulation for Uniform Acceleration? Acceleration: a(t) = 292 Velocity: v'(t) = a(t), thus v(t) = 292t + 1 Position: x'(t) = v(t), thus x(t) = 146t^2 + t - 24 We didn't really need x(t) for this problem. What is v(1.1)?

OpenStudy (anonymous):

I just solved it, v2=v1+acceleration(time) 1m/s+ 292(1.1) = 322.2 m/s

OpenStudy (aum):

Correct. v = u + at = 1 + 292(1.1) = 322.2 m/s

OpenStudy (tkhunny):

a(1.1) = 292*1.1 + 1 = 322.2 -- There you go.